OpenStudy (anonymous):

40% of what number is 100?

OpenStudy (anonymous):

250

OpenStudy (md152727):

Yep, 250, http://www.percentagecalculator.net/

OpenStudy (anonymous):

Wow, how exactly did you get that answer?

OpenStudy (anonymous):

I'll just do a cross multiply and I should get the answer.

OpenStudy (md152727):

on the second bar, write 100 is what percent of 40? then click calucate

OpenStudy (md152727):

please medal

OpenStudy (unklerhaukus):

\[40\% \times n = 100\\ \frac{40}{100}\times n = 100\\ n = 100\times\frac{100}{40}\\ \qquad=\frac{10000}{40}\\ \qquad=\frac{1000}4\\ \qquad=\]